OpenStudy (anonymous):

A man has walked 30% of his journey and finds that he has only walked 27km. how long is his whole journey?

OpenStudy (anonymous):

divide 27 by 0.3

OpenStudy (anonymous):

divide 27 by 3 and multiply the result by ten. so 27/3=9 then 9*10=90.

OpenStudy (lilg132):

27 * 100 = 2700 2700 / 30 = 90km

OpenStudy (lilg132):

dont forget to click good answer :)

OpenStudy (anonymous):

let the journey 2 walk be x, then 30% of x=27 or 0.3*x=27 x=27/0.3 so, x=90km ans